# Adaptive Exploration for Multi-Reward Multi-Policy Evaluation

### v2

@@ Abstract
We study the policy evaluation problem in an online multi-reward multi-policy discounted setting, where multiple reward functions must be evaluated simultaneously for different policies. We adopt an $(\epsilon,\delta)$-PAC perspective to achieve $\epsilon$-accurate estimates with high confidence across finite or convex sets of rewards, a setting that has not been investigated in the literature. Building on prior work on Multi-Reward Best Policy Identification, we adapt the MR-NaS exploration scheme (Russo & Vannella, 2024) to jointly minimize sample complexity for evaluating different policies across different reward sets. Our approach leverages an instance-specific lower bound revealing how the sample complexity scales with a measure of value deviation, guiding the design of an efficient exploration policy. Although computing this bound entails a hard non-convex optimization, we propose an efficient convex approximation that holds for both finite and convex reward sets. Experiments in tabular domains demonstrate the effectiveness of this adaptive exploration scheme [^1] Code repository: https://github.com/rssalessio/multi-reward-multi-policy-eval..
@@ 1 Introduction
Our result exhibits a sample complexity that scales with the worst-case policy-reward pair—rather than the sum over individual policy-reward pairs—capturing the inherent difficulty of efficiently evaluating all combinations simultaneously. Particularly, the complexity scales according to a measure of value deviation $\rho_{r}^{\pi}(s,s^{\prime})\coloneqq V_{r}^{\pi}(s^{\prime})-\mathbb{E}_{\hat{s}\sim P(s,\pi(s))}[V_{r}^{\pi}(\hat{s})]$, similar to the variance of the value:
@@ 3 Problem Setting
In this section, we describe the MDP model considered, and the policy evaluation setting.
@@ 3.1 Markov Decision Processes (MDPs)
Markov Decision Processes (MDPs) are widely utilized to model sequential decision-making tasks (Puterman, 2014). In these tasks, at each time-step $t\in\mathbb{N}$ an agent observes the current state of the MDP $s_{t}$ and selects an action $a_{t}$ to achieve a desired objective. This objective is encapsulated in terms of a reward $r_{t}\in[0,1]$, observed after selecting the action. In RL, the primary goal of the agent is to determine a sequence of actions that maximizes the total reward collected from an unknown MDP.
@@
We consider a discounted MDP (Puterman, 2014) of the type $M=({\cal S},{\cal A},P,\gamma)$, where: ${\cal S}$ is a finite state space of size $S=|{\cal S}|$; ${\cal A}$ is a finite action space of size $A=|{\cal A}|$; $P:{\cal S}\times{\cal A}\to\Delta({\cal S})$ is the transition function, which maps state-action pairs to distributions over states; lastly, $\gamma\in(0,1)$ is the discount factor. In the following we consider also reward functions $r:{\cal S}\times{\cal A}\to[0,1]$ that are bounded and deterministic functions of state-action pairs, and write $M_{r}=(M,r)$ to denote the MDP $M$ with reward $r$. In classical RL an agent is interested in maximizing the total discounted reward collected from $M_{r}$: $r_{1}+\gamma r_{2}+\gamma^{2}r_{3}+\dots$. The problem of computing an optimal sequence of actions (i.e., that maximizes the total discounted reward) can be reduced to that of finding a Markovian policy $\pi:S\to\Delta({\cal A})$ that maps states to distributions over actions (Puterman, 2014). For a Markovian policy $\pi$ we define the discounted value of $\pi$ in $M_{r}$ at state $s$ as $V_{r}^{\pi}(s)\coloneqq\mathbb{E}^{\pi}[\sum_{t\geq 1}\gamma^{t-1}r(s_{t},a_{t})|s_{1}=s]$, where $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ and $a_{t}\sim\pi(\cdot|s_{t})$ (if $\pi$ is deterministic, we write $a_{t}=\pi(s_{t})$). We also write $V_{r}^{\pi}$, and omit the dependency on $M$ when it is clear from the context. We also define $V_{r}^{\star}(s)\coloneqq\max_{\pi}V_{r}^{\pi}(s)$ to be the optimal value in $s$ over all Markovian policies. We further define the action-value function in $M_{r}$ as $Q_{r}^{\pi}(s,a)\coloneqq r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V_{r}^{\pi}(s^{\prime})]$.
Lastly, for two MDPs $M,M^{\prime}$ with transition functions, respectively, $P,P^{\prime}$, we write ${\rm KL}_{P|P^{\prime}}(s,a)$ to indicate the KL-divergence between the two transition functions in $(s,a)\in{\cal S}\times{\cal A}$. We also write ${\rm kl}(x,y)=x\log(x/y)+(1-x)\log((1-x)/(1-y))$ to denote the Bernoulli KL-divergence between two parameters $x,y\in[0,1]$.
@@
We consider the problem of evaluating a finite set of deterministic target policies $\Pi=\{\pi_{1},\dots,\pi_{N}\}$ over finite, or possibly convex, reward sets (we discuss more on this in the next paragraph). On the MDP $M$ we impose the following assumption.
@@
The MDP $M$ is communicating, and the learner has no prior knowledge of $P$.
As in (Al Marjani et al., 2021; Russo & Vannella, 2024), we assume the MDP to be communicating, which avoids the awkward case where the algorithm could enter a subclass of states from which there is no possible comeback, and thus it becomes impossible to identify the value of a policy $\pi$ to the desired accuracy. While the results can be extended to weakly-communicating MDPs, for ease of exposition we leave the extension for future work.
@@
In the following we consider sets of reward functions ${\cal R}$ that are either finite, or convex, and assume that these sets are known to the agent beforehand. For tabular MDPs (i.e., with finite state-action spaces), assuming a natural labeling of the states ${\cal S}=\{s_{1},\dots,s_{S}\}$, we represent a reward $r$, for a policy $\pi$, as a vector in $[0,1]^{S}$ (for simplicity, we omit the dependency on $\pi$). The $i$-th element of $r$ corresponds to the reward $r(s_{i},\pi(s_{i}))$. For $\pi$, we indicate the canonical reward set by ${\cal R}_{\rm canon}^{\pi}=\{e_{1},\dots,e_{S}\}$ in $\mathbb{R}^{S}$, with $e_{i}$ being the $i$-th vector of the canonical basis, defined as $(e_{i})_{j}=\mathbf{1}_{(i=j)},\;i,j\in\{1,\dots,S\}$. Lastly, we denote the ${\cal R}=[0,1]^{S}$ case as reward-free since it encompasses all possible reward for $\pi$.
@@ 3.2 Online Multi-Reward Multi-Policy Evaluation
This setting is closely related to off-policy policy evaluation (OPE) (Thomas & Brunskill, 2016; Precup, 2000). However, OPE does not typically focuses on the problem of optimizing the data collection policy. In this work, we study the problem of devising an optimal data-collection policy that, by online interactions with the MDP $M$, permits the agent to learn the value of pairs $(\pi,{\cal R}_{\pi})\in\Theta$ as quickly as possible up to the desired accuracy.
@@
  • an estimated value $\hat{V}_{r}$: at the stopping time $\tau$, Alg returns the estimated value $\hat{V}_{r}^{\pi}$ of the policy $\pi\in\Pi$ in $M$ for any chosen reward $r\in{\cal R}_{\pi}$.
@@
An algorithm Alg is said to be multi-reward multi-policy $(\epsilon,\delta)$-PAC if, for any MDP $M$, policies-rewards set $\Theta$, $\delta\in(0,1/2),\epsilon\in(0,\frac{1}{2(1-\gamma)})$, we have $\mathbb{P}_{M}[\tau<\infty]=1$ (the algorithm stops almost surely) and
@@
In the following section, we investigate the sample complexity of this problem and determine the minimal number of samples required to achieve the $(\epsilon,\delta)$-PAC guarantees. Our analysis reveals that the problem can be more challenging under certain rewards. In fact, the sample complexity is governed by the most difficult reward-policy pair in $\Theta$, rather than a sum of the individual complexities across pairs.
@@ 4.1 Set of Alternative Models
Confusing models are alternative models that are “similar” to $M$, but differ in certain key properties. As we see later, the set of alternative models is crucial for establishing a lower bound on the sample complexity $\tau$. To prove the lower bound, we frame the sample complexity problem as a goodness-of-fit test: does the observed data better align the true model $M$ or an alternative model $M^{\prime}$?
@@
where $M_{r}^{\prime}=({\cal S},{\cal A},P_{r}^{\prime},r,\gamma)$ is an alternative MDP, and we index the transition function $P_{r}^{\prime}$ by the corresponding reward $r$ for clarity. The notation $M\ll M_{r}^{\prime}$ means that $P(s,a)$ is absolutely continuous with respect to $P_{r}^{\prime}(s,a)$ for all $(s,a)$ (and $P$ is the transition function of $M$). We also denote by ${\rm Alt}^{\epsilon}(M)=\cup_{\pi\in\Pi,r\in{\cal R}_{\pi}}{\rm Alt}_{\pi,r}^{\epsilon}(M)$ the entire set of alternative models over ${\cal R}$.
@@
Note that we require $\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>2\epsilon$, rather than just $\epsilon$. If the separation were only $\epsilon$, an estimator $\hat{V}_{r}^{\pi}$ satisfying $\|\hat{V}_{r}^{\pi}-V_{M_{r}}^{\pi}\|_{\infty}\leq\epsilon$ could also be simultaneously $\epsilon$-accurate for an alternative model $M^{\prime}$ satisfying $\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>\epsilon$, making the models indistinguishable. Requiring $2\epsilon$ separation prevents this ambiguity. While this choice potentially weakens the resulting lower bound, we believe the looseness is at most a constant factor (cf.upper bound in Theorem 5.2).
Lastly, consider the case where the set ${\rm Alt}_{\pi,r}^{\epsilon}(M)$ is empty for some pair $(\pi,r)$. In this situation, any model $P^{\prime}$ sharing the same support as $P$ yields a value that is $2\epsilon$-close to the true value. Thus, the learning challenge for that reward is rather minimal. Therefore, characterizing when these sets are empty is crucial in the analysis.
@@
This quantity measures how much the value at a state $s^{\prime}$ deviates from the expected value under $\pi$ at $s$. As we see later it is “easier” to construct alternative models if $|\rho_{r}^{\pi}(s,s^{\prime})|$ is large.
We also define these quantities in vector form $\rho_{r}^{\pi}(s)\coloneqq\begin{bmatrix}\rho_{r}^{\pi}(s,s_{1})&\dots&\rho_{r}^{\pi}(s,s_{S})\end{bmatrix}^{\top}$, so that $\|\rho_{r}^{\pi}(s)\|_{\infty}=\max_{s^{\prime}}|\rho_{r}^{\pi}(s,s^{\prime})|$ is the maximum one-step deviation at $s$. The deviation $\rho_{r}^{\pi}$ is closely related to the span of the value function ${\rm sp}(V_{r}^{\pi})\coloneqq\max_{s^{\prime}}V_{r}^{\pi}(s^{\prime})-\min_{s^{\prime}}V_{r}^{\pi}(s)$, but, is in general smaller, as discussed in Section B.1.2.
@@
Fix a policy $\pi\in\Pi$ and a reward $r\in\mathcal{R}_{\pi}$. We have the following conditions: (1; necessary) if $\forall s,\|\rho_{r}^{\pi}(s)\|_{\infty}\leq\frac{\epsilon(1-\gamma)}{\gamma}$, then ${\rm Alt}_{\pi,r}^{\epsilon}(M)=\emptyset$; (2; sufficient) if $\exists s\in{\cal S}:\|\rho_{r}^{\pi}(s)\|_{\infty}>\frac{2\epsilon}{\gamma}$, then ${\rm Alt}_{\pi,r}^{\epsilon}(M)\neq\emptyset$.
@@
  • As $\epsilon$ increases or $\gamma$ decreases, the necessary condition in Proposition 4.1 is more easily satisfied, potentially leading to a decrease in the number of alternative models. Consequently, this may imply a lower sample complexity.
  • The proof highlights the concept of confusing states: a state $s$ is considered confusing if $\|\rho_{r}^{\pi}(s)\|_{\infty}>\frac{2\epsilon}{\gamma}$. States with smaller values of $\|\rho_{r}^{\pi}(s)\|_{\infty}$ do not strongly affect the sample complexity. Additionally, Lemma B.3 (in the appendix) indicates that the maximum $\max_{s^{\prime}}|\rho_{r}^{\pi}(s,s^{\prime})|$ is typically attained at $s\neq s^{\prime}$.
  • Since $|\rho_{r}^{\pi}(s,s^{\prime})|\leq{\rm sp}(V_{r}^{\pi})$ (as shown in Lemma B.3 in the appendix), similar values across states (smaller span) reduce the potential for confusing models, thereby lowering the sample complexity.
@@
The vectors $r$ for which $\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}=0$ is precisely the subspace $\{\alpha{\bf 1}:\alpha\in[0,1]\}$, where ${\bf 1}$ is the vector of ones.
While it may seem obvious that the unit vector reward cannot produce an alternative model, it is noteworthy that no other reward satisfies $\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}=0$ for any MDP.

### v3

@@ Abstract
We study the policy evaluation problem in an online multi-reward multi-policy discounted setting, where multiple reward functions must be evaluated simultaneously for different policies. We adopt an $(\epsilon,\delta)$-PAC perspective to achieve $\epsilon$-accurate estimates with high confidence over finite or convex sets of rewards, a setting that has not been systematically studied in the literature. Building on prior work on Multi-Reward Best Policy Identification, we adapt the MR-NaS exploration scheme (Russo & Vannella, 2024) to jointly minimize sample complexity for evaluating different policies across different reward sets. Our approach leverages an instance-specific lower bound revealing how the sample complexity scales with a measure of value deviation, guiding the design of an efficient exploration policy. Although computing this bound entails a hard non-convex optimization, we propose an efficient convex approximation that holds for both finite and convex reward sets. Experiments in tabular domains demonstrate the effectiveness of this adaptive exploration scheme. Code repository: https://github.com/rssalessio/multi-reward-multi-policy-eval.
@@ 1 Introduction
[…]
@@ 3 Problem Setting
In this section, we describe the MDP model considered and the policy evaluation setting.
@@ 3.1 Markov Decision Processes (MDPs)
Markov Decision Processes (MDPs) are widely utilized to model sequential decision-making tasks (Puterman, 2014). In these tasks, at each time-step $t=1,2,\dots$ an agent observes the current state of the MDP $s_{t}$ and selects an action $a_{t}$ to achieve a desired objective. This objective is encapsulated in terms of a reward $r_{t}\in[0,1]$, observed after selecting the action. In RL, the primary goal of the agent is to determine a sequence of actions that maximizes the total reward collected from an unknown MDP.
@@
We consider a discounted MDP (Puterman, 2014) of the type $M=({\cal S},{\cal A},P,\gamma)$, where: $\cal S$} is a finite state space of size $S=|{\cal S}|$; $\cal A$} is a finite action space of size $A=|{\cal A}|$; $P:{\cal S}\times{\cal A}\to\Delta({\cal S})$ is the transition function, which maps state-action pairs to distributions over states; lastly, $\gamma\in(0,1)$ is the discount factor. In the following we consider also reward functions $r:{\cal S}\times{\cal A}\to[0,1]$ that are bounded and deterministic functions of state-action pairs, and write $M_{r}=(M,r)$ to denote the MDP $M$ with reward $r$. In classical RL, an agent is interested in maximizing the total discounted reward collected from $M_{r}$: $r_{1}+\gamma r_{2}+\gamma^{2}r_{3}+\dots$. The problem of computing an optimal sequence of actions (i.e., that maximizes the total discounted reward) can be reduced to that of finding a Markovian policy $\pi:{\cal S}\to\Delta({\cal A})$ that maps states to distributions over actions (Puterman, 2014). For a Markovian policy $\pi$ we define the discounted value of $\pi$ in $M_{r}$ at state $s$ as $V_{r}^{\pi}(s)\coloneqq\mathbb{E}^{\pi}[\sum_{t\geq 1}\gamma^{t-1}r(s_{t},a_{t})|s_{1}=s]$, where $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ and $a_{t}\sim\pi(\cdot|s_{t})$ (if $\pi$ is deterministic, we write $a_{t}=\pi(s_{t})$). We also write $V_{r}^{\pi}$ and omit the dependency on $M$ when it is clear from the context. We also define $V_{r}^{\star}(s)\coloneqq\max_{\pi}V_{r}^{\pi}(s)$ to be the optimal value in $s$ over all Markovian policies. We further define the action-value function in $M_{r}$ as $Q_{r}^{\pi}(s,a)\coloneqq r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V_{r}^{\pi}(s^{\prime})]$.
Additional notation.
For a set ${\cal U}\subset\mathbb{R}^{n},n\in\mathbb{N},$ we denote by $\overline{\cal U}$ its closure, and we also define $\Delta({\cal U})$ to be the set of distributions over $\cal U$}. We also denote by $N_{t}(s,a)=\sum_{n=1}^{t}\mathbf{1}_{((s_{n},a_{n})=(s,a))}$ the number of times an algorithm has visited a state-action pair $(s,a)$ up to time $t$ (sim. we define $N_{t}(s)=\sum_{a}N_{t}(s,a)$). For two MDPs $M,M^{\prime}$ with transition functions, respectively, $P,P^{\prime}$, we write ${\rm KL}_{P|P^{\prime}}(s,a)$ to indicate the KL-divergence between the two transition functions in $(s,a)\in{\cal S}\times{\cal A}$ (we also write ${\rm KL}_{M|M^{\prime}}(s,a)$). We denote by ${\rm kl}(x,y)=x\log(x/y)+(1-x)\log((1-x)/(1-y))$ the Bernoulli KL-divergence between two parameters $x,y\in[0,1]$.
@@
We consider the problem of evaluating a finite set of deterministic target policies $\Pi=\{\pi_{1},\dots,\pi_{N}\}$ over finite, or possibly convex, reward sets (we discuss more on this in the next paragraph). On the MDP $M$, we impose the following assumption.
@@
The MDP $M$ is communicating and aperiodic under a uniform policy, and the learner has no prior knowledge of $P$. The starting state is also arbitrary.
As in (Al Marjani et al., 2021; Russo & Vannella, 2024), we assume that the MDP is communicating, which avoids the awkward case where the algorithm could enter a subclass of states from which there is no possible comeback, and thus it becomes impossible to identify the value of a policy $\pi$ to the desired accuracy. Lastly, the assumption of aperiodicity can be met by assuming that the exploration policy assigns positive probability to some state-action pair $(s,a)$ satisfying $P(s|s,a)>0$. Such assumption is needed to ensure ergodicity of the chain induced by the exploration policy.
@@
[…]
@@ 3.2 Online Multi-Reward Multi-Policy Evaluation
This setting is closely related to off-policy policy evaluation (OPE) (Thomas & Brunskill, 2016; Precup, 2000). However, OPE does not typically focus on the problem of optimizing the data collection policy. In this work, we study the problem of devising an optimal data-collection policy that, by online interactions with the MDP $M$, permits the agent to learn the value of pairs $(\pi,{\cal R}_{\pi})\in\Theta$ as quickly as possible up to the desired accuracy.
@@
  • an estimated value $\hat{V}_{r}^{\pi}$: at the stopping time $\tau$, Alg returns the estimated value $\hat{V}_{r}^{\pi}$ of the policy $\pi\in\Pi$ in $M$ for any chosen reward $r\in{\cal R}_{\pi}$.
@@
An algorithm Alg is said to be multi-reward multi-policy $(\epsilon,\delta)$-PAC if, for any MDP $M$, policies-rewards set $\Theta$, $\delta\in(0,1/2)$ and $\epsilon\in(0,\frac{1}{2(1-\gamma)})$, we have $\mathbb{P}_{M}[\tau<\infty]=1$ (the algorithm stops almost surely) and
@@
In the following section, we investigate the sample complexity of this problem and determine the minimal number of samples required to achieve the $(\epsilon,\delta)$-PAC guarantees. Our analysis reveals that the problem can be more challenging under certain rewards. In fact, the sample complexity is governed by the most difficult policy-reward pair in $\Theta$, rather than the sum of the individual complexities across pairs.
@@ 4.1 Set of Alternative Models
Confusing models are alternative models that are “similar” to $M$, but differ in certain key properties. As we see later, the set of alternative models is crucial for establishing a lower bound on the sample complexity $\tau$. To prove the lower bound, we frame the sample complexity problem as a goodness-of-fit test: does the observed data better align with the true model $M$ or an alternative model $M^{\prime}$?
@@
[…]
@@
Note that we require $\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>2\epsilon$, rather than just $\epsilon$. If the separation were only $\epsilon$, an estimator $\hat{V}_{r}^{\pi}$ satisfying $\|\hat{V}_{r}^{\pi}-V_{M_{r}}^{\pi}\|_{\infty}\leq\epsilon$ could also be simultaneously $\epsilon$-accurate for an alternative model $M^{\prime}$ satisfying $\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>\epsilon$, making the models indistinguishable. Requiring $2\epsilon$ separation prevents this ambiguity. While this choice potentially weakens the resulting lower bound, we believe the looseness is at most a constant factor (see Theorem 5.2).
Lastly, consider the case where the set ${\rm Alt}_{\pi,r}^{\epsilon}(M)$ is empty for some pair $(\pi,r)$. In this situation, any model $P^{\prime}$ sharing the same support as $P$ yields a value that is $2\epsilon$-close to the true value. Thus, the learning challenge for that reward is rather minimal. Therefore, characterizing when these sets are empty is crucial in the analysis. In the following, we denote by ${\cal R}_{\pi}^{\epsilon}=\{r\in{\cal R}_{\pi}:{\rm Alt}_{\pi,r}^{\epsilon}(M)\neq\emptyset\}$ the set of rewards for which the corresponding set of confusing models is non-empty.
@@
This quantity measures how much the value at a state $s^{\prime}$ deviates from the expected value under $\pi$ at $s$. As we see later, it is “easier” to construct alternative models if $|\rho_{r}^{\pi}(s,s^{\prime})|$ is large. We also define these quantities in vector form $\rho_{r}^{\pi}(s)\coloneqq\begin{bmatrix}\rho_{r}^{\pi}(s,s_{1})&\dots&\rho_{r}^{\pi}(s,s_{S})\end{bmatrix}^{\top}$, so that $\|\rho_{r}^{\pi}(s)\|_{\infty}=\max_{s^{\prime}}|\rho_{r}^{\pi}(s,s^{\prime})|$ is the maximum one-step deviation at $s$. The deviation $\rho_{r}^{\pi}$ is closely related to the span of the value function ${\rm sp}(V_{r}^{\pi})\coloneqq\max_{s^{\prime}}V_{r}^{\pi}(s^{\prime})-\min_{s}V_{r}^{\pi}(s)$, but, is in general smaller than the span (see Lemma B.3).
@@
Fix a policy $\pi\in\Pi$ and a reward $r\in\mathcal{R}_{\pi}$. We have the following conditions: (1; necessary) if ${\rm Alt}_{\pi,r}^{\epsilon}(M)\neq\emptyset$ then $\exists s\in{\cal S}:\|\rho_{r}^{\pi}(s)\|_{\infty}>\frac{\epsilon(1-\gamma)}{\gamma}$; (2; sufficient) if $\exists s\in{\cal S}:\|\rho_{r}^{\pi}(s)\|_{\infty}>\frac{2\epsilon}{\gamma}$, then ${\rm Alt}_{\pi,r}^{\epsilon}(M)\neq\emptyset$.
@@
  • As $\epsilon$ increases or $\gamma$ decreases, the necessary condition in Proposition 4.1 is less likely to be satisfied, potentially leading to a decrease in the number of alternative models (which suggests a smaller sample complexity).
  • The proof highlights the concept of confusing states: a state $s$ is considered confusing if $\|\rho_{r}^{\pi}(s)\|_{\infty}>\frac{2\epsilon}{\gamma}$. States with smaller values of $\|\rho_{r}^{\pi}(s)\|_{\infty}$ do not strongly affect the sample complexity. Additionally, Lemma B.3 (in the appendix) indicates that the maximum $\max_{s^{\prime}}|\rho_{r}^{\pi}(s,s^{\prime})|$ is typically attained at $s^{\prime}\neq s$.
@@
The vectors $r$ for which $\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}=0$ is precisely the set $\{\alpha{\bf 1}:\alpha\in[0,1]\}$, where $\bf 1$} is the vector of ones.
While it may seem obvious that the all-ones reward cannot produce an alternative model, it is noteworthy that no other reward satisfies $\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}=0$.